ON THE SECOND MOMENT FOR PRIMES IN AN ARITHMETIC PROGRESSION

DANIEL GOLDSTON C. Y. YILDIRIM

Abstract. Assuming the Generalized Riemann Hypothesis, we obtain a lower bound within a constant factor of the conjectured asymptotic result for the second moment for primes in an individual arithmetic progression in short intervals. Previous results were averaged over all progression of a given modulus. The method uses a short divisor sum approximation for the von Mangoldt function, together with some new results for binary correlations of this divisor sum approximation in arithmetic progressions.